OpenStudy (anonymous):

a bag contains 3 red marbles, 4 green marbles.. A marble is drawn from a bag and not put back before 2nd marble is drawn. What is the probability that 2nd marble drawn is red?

OpenStudy (anonymous):

So either the first marble can be red AND the second marble can be red, or the first marble can be geen and the second marble can be red. What do you think you have to do?

OpenStudy (anonymous):

P(E) = (3/7 * 2/6) + (4/7 * 3/6) = ....

OpenStudy (anonymous):

Also, if you were wondering, there are 7 marbles in the bag, but when you take one out, you do not put it back, so thre's only 6 left.

OpenStudy (anonymous):

the probability is first marble red,second red or first green then second red i.e. P(RR) OR P(GR) \[\frac{ 3 }{ 7 }\times \frac{ 2 }{ 6 }+\frac{ 4 }{ 7 }\times \frac{ 3 }{ 6 }=\frac{ 1 }{ 7 }+\frac{ 2 }{ 7}=\frac{ 3 }{ 7 }\]

OpenStudy (anonymous):

ready to get help?

OpenStudy (anonymous):

So, we are starting with 3 red, 4 green, 7 total marbles. And we draw one out. Lets start with the first one.

OpenStudy (anonymous):

It could be red, or green, right?

OpenStudy (anonymous):

@nylejre are you there?

OpenStudy (anonymous):

Do you know how to find the probability that it could be green?

OpenStudy (anonymous):

if not. When looking at just drawing one of the marbles out, we can find the probability of a certain color by dividing the number of marbles of that color by the total marbles in the bag.

OpenStudy (anonymous):

I'll have to leave sine your not ready to get help.

OpenStudy (anonymous):

I have 3/7 because there are 3 possible red marbles to draw, and 7 total marbles.

OpenStudy (anonymous):

so do the same with the green 4/7

OpenStudy (anonymous):

Now, we know that the first marble chosen is not replaced, so the number of marbles in the bag is now 6, correct?

OpenStudy (anonymous):

we can draw this if you like, just depends.

OpenStudy (anonymous):

next we want to find the probability that the second marble drawn is also red.

OpenStudy (anonymous):

So. We have 2 different options. We could have the 1st marble be red, and the 2nd marble be red. Or, we could have the 1st marble be green, and the 2nd marble be red.

OpenStudy (anonymous):

Lets look at option 1, both marbles are red.

OpenStudy (anonymous):

so what's the probability that the 2nd marble will be red?

OpenStudy (anonymous):

Note* It could be either one, you have to consider both options, unless you are told otherwise.

OpenStudy (anonymous):

Ok. So now we have 6 total marbles, and only 2 red

OpenStudy (anonymous):

ok wait.. i'll just analyzed the first thing you did..

OpenStudy (anonymous):

Since we have 2 red left, and 6 marbles, the probability will be 2/6. Now that we know that, we need to find the probabilty that both draws are red.

OpenStudy (anonymous):

alright then. I'll continue with me explanation in the meanwhile.

OpenStudy (anonymous):

Since we have 2 red left, and 6 marbles, the probability will be 2/6. Now that we know that, we need to find the probabilty that both draws are red. To do that, we multiply our probability that the first draw is red (3/7) by the probability that the 2nd draw is also red (2/6)

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

Now we can work on the second option, 1 green, 1 red.

OpenStudy (anonymous):

To multiply fractions, you can use your calculator to get a decimal, or you can multiply the numerators by the denominators... For example, (3*2)/(7*6) That will give you a fraction for your answer. Either should be fine

OpenStudy (anonymous):

1/7

OpenStudy (anonymous):

Yup good!

OpenStudy (anonymous):

are you clear up to that?

OpenStudy (anonymous):

Okay that was a part, now we have to find the probability of the second option, which is one green, 1 red.

OpenStudy (anonymous):

Now, we find the probability that the 2nd marble is red.

OpenStudy (anonymous):

If the first green marble isn't replaced, we have only 6 marbles, and 3 of them are red, correct? so...

OpenStudy (anonymous):

... Now, we have mulitply these probabilities to find the probability that the first marble drawn is green AND the second marble drawn is red.

OpenStudy (anonymous):

4/7 * 3/6?

OpenStudy (anonymous):

2/7

OpenStudy (anonymous):

= 2/7 correct? :)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

any questions on that?

OpenStudy (anonymous):

Ok. For the second marble to be red, we could have either one of these 2 options occur, so the total probability that the 2nd marble is red is the sum of these 2 probabilities.

OpenStudy (anonymous):

or 1/7 + 2/7

OpenStudy (anonymous):

what is 1/7+ 2/7?

OpenStudy (anonymous):

the answer is 3/7 Yes. That should be your final answer.

OpenStudy (anonymous):

I hope I help answer your question, and hopefully by now your confusion is recovered. Have a great day, and feel free to ask any other questions.

OpenStudy (anonymous):

Bye for now!

OpenStudy (anonymous):

waahh... you did great!! I understand it better now.. thanks a lot!

OpenStudy (anonymous):

anything else before I leave, are you all confirmed?

OpenStudy (anonymous):

That is great to hear, I hope your feeling much better with these kind of problems.